THE ν_{5} AND 2ν_{9} BANDS OF THE^{15}N ISOTOPIC SPECIES OF NITRIC ACID (H^{15}NO_{3}): LINE POSITIONS AND INTENSITIES

\underline{A. PERRIN}, Laboratoire Inter Universitaire des Systemes Atmosphériques, CNRS, Université Paris 12, 61 Av du Général de Gaulle, 94010 Créteil Cedex France; R. MBIAKE, Centre de Physique Atomique et Moléculaire et Optique Quantique (CEPAMOQ), Université de Douala, B.P. 8580, Cameroun.

We present the first high resolution Fourier transform analysis of the 11 μm bands for H^{15}NO_{3} which is the second most abundant isotopomer of nitric acid (a=0.00365). In this way the analysis of the ν_{5} and 2ν_{9} cold bands centered at 871.095 and 893.452 cm^{−1} was performed. As for H^{14}NO_{3}, these bands are significantly perturbed since rather strong resonances couple the 5^{1} and 9^{2} rotational levels. The theoretical model that we used to compute the line positions and line intensities is directly issued from the one which we used recently for H^{14}NO_{3} ^{a}. Actually for the H^{15}NO_{3} line positions, the Hamiltonian matrix accounts for the rather strong Fermi and the weaker Coriolis interactions linking the v=5^{1} and v=9^{2} rotational energy levels. Using this model which describes correctly the strong mixing of the 5^{1} and 9^{2} upper state energy levels, the ν_{5} and 2ν_{9} line intensities for H^{15}NO_{3} were satisfactorily computed using the ν_{5} and 2ν_{9} transition moment operators achieved previously for the ^{14}N (main) isotopic species. In this way, the transfer of intensities from the ν_{5} fundamental (and presumably strong) band to the 2ν_{9} overtone (and presumably weak) band could be explained for H^{15}NO_{3} as it was done previously for the ^{14}N (main) isotopic species.

aA.Perrin, J.Orphal, J.-M.Flaud, S.Klee, G.Mellau, H.Maeder, D.Walbrodt, and M.Winnewisser, (2004), J. Mol. Spectrosc. 228, 375-391